I have no idea how to even begin to solve this problem. Help please!
An automobile license plate consists of 3 letters followed by 4 digits.

How many different plates can be made?
If repetitions are allowed?
If repetitions are not allowed?
If repetitions are allowed in the letters but not in the digits?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

I have no idea how to even begin to solve this problem. Help please!
An automobile license plate consists of 3 letters followed by 4 digits.
How many different plates can be made?
If repetitions are allowed?
Assuming it's all CAPs, you get
26*26*26 for the letters, then
10*10*10*10 for the numbers.
= 26*26*26*10*10*10*10 = 175760000.
------------------
If repetitions are not allowed?
The 1st letter is 1 of 26, then 1 of 25, then 24.
Then 1 of 10 numbers, 1 0f 9, etc.
= 26*25*24*10*9*8*7 = 78624000.
-----------------
If repetitions are allowed in the letters but not in the digits?
Do it like the ones above.

Answer by stanbon(75887)

An automobile license plate consists of 3 letters followed by 4 digits.
How many different plates can be made?
If repetitions are allowed?-----26^3*10^4
If repetitions are not allowed?---26*25*24*10*9*8 =
If repetitions are allowed
in the letters but not in the digits?---26^3*10*9*8
====================================================================
Cheers,
Stan H.